OpenStudy (rainbow_dash):

Find the sum of the given infinite geometric series. 3/2 - 3/4 + 3/8 ...

OpenStudy (anonymous):

\[\Huge \frac{3}{2} - \frac{3}{4} + \frac{3}{8} - \frac{3}{16} ... = \] \[\Huge \sum_{n = 1}^{\infty} -\frac{3}{(-2)^{n}} = -3\sum_{n=1}^{\infty}\frac{1}{(-2)^n}\]

OpenStudy (rainbow_dash):

i don't really understand sigma all that well :(